OpenStudy (anonymous):

Given the function f(x) = 2x, find the value of f−1(16)

OpenStudy (anonymous):

say, f(x)=y so y=2x to find the inverse function, always interchange x and y so the inverse function is 2y=x or, y=x/2 so f^-1 (x)=x/2 so f-1 (16)=16/2=8

OpenStudy (anonymous):

Thank you :) could you help with another?

OpenStudy (anonymous):

please post a separate question. thank you!